### v1

# Scalable DC Optimization via Adaptive Frank-Wolfe Algorithms

Abstract
We consider the problem of minimizing a difference of (smooth) convex functions over a compact convex feasible region $P$, i.e., $\min_{x\in P}f(x)-g(x)$, with smooth $f$ and Lipschitz continuous $g$. This computational study builds upon and complements the framework of Maskan et al. [2025] by integrating advanced Frank-Wolfe variants to reduce computational overhead. We empirically show that constrained DC problems can be efficiently solved using a combination of the Blended Pairwise Conditional Gradients (BPCG) algorithm [Tsuji et al., 2022] with warm-starting and the adaptive error bound from Maskan et al. [2025]. The result is a highly efficient and scalable projection-free algorithm for constrained DC optimization.
1 Introduction
In this paper, we focus on solving constrained Difference-of-convex (DC) optimization of the form
$\min_{x\in P}\Phi(x)\doteq\min_{x\in P}f(x)-g(x),$ (1.1)
where $P$ is a compact convex feasible region, $f$ is an $L_{f}$-smooth convex function, and $g$ is a $L_{g}$-Lipschitz continuous convex function, following the setup in Maskan et al. [2025]. This problem falls into the class of DC programming problems, which are typically nonconvex but still somewhat well-structured optimization problems. We are in particular interested in the case where the feasible region $P$ is complicated, so that traditional projection-based methods are computationally prohibitive within the DCA algorithm and projection-free methods, in particular the Frank-Wolfe algorithm, are a natural choice given that they only rely on a linear minimization oracle to access $P$, which is often much cheaper than projections [Combettes & Pokutta, 2021]. We require access to the gradients of $f$, subgradients of $g$ via (sub-)gradient oracles and access to $P$ via a linear minimization oracle.
Related Work
Both DC Programming as well as Frank-Wolfe methods are very active fields of research. Here we will only be able to briefly discuss works that are most relevant to our setting here.
DC Programming
The difference-of-convex optimization framework provides a powerful mathematical structure for tackling a broad class of structured nonconvex optimization problems. The fundamental algorithmic approach for this problem class is the Difference-of-Convex Algorithm (DCA), first developed by Tao et al. [1986] and over the last thirty years DCA and the DC framework have established themselves as particularly relevant in practical context, such as, e.g., indefinite quadratic optimization [An & Tao, 1997, 1998], convex-concave minimax problems [Yuille & Rangarajan, 2003, Shen et al., 2016], machine learning kernel selection [Argyriou et al., 2006], hierarchical bilevel optimization [Hoai An et al., 2009], contextual decision making under uncertainty [Bennouna et al., 2024], domain adaptation with distribution discrepancy [Awasthi et al., 2024a], structured sampling via determinantal processes [Mariet & Sra, 2015], adversarial robustness in deep learning [Awasthi et al., 2024b], and more recently in generalized polyhedral DC problems Huong et al. [2024]. An efficient practical improvement to the DCA algorithm is boosted DCA, where the new iterate of the DCA does not simply replace the old one but a line search is performed, which can improve computational performance in many cases [Aragón Artacho & Vuong, 2018, Aragón Artacho et al., 2018, Aragón Artacho et al., 2022, Ferreira et al., 2022]. This improvement via an additional line search can also be readily combined with the approach presented here, however in initial experiments we did not find significant additional improvements over the speed-up offered by BPCG with warm-starting and adaptive error bounds (see Section B). An in-depth analysis, which is beyond the scope of this paper, is left for follow-up work. Also the theoretical understanding of DCA has evolved significantly over time. Initially, Tao [1997] established the asymptotic convergence properties, which were subsequently refined by Lanckriet & Sriperumbudur [2009] through simpler analyses under differentiability conditions; see also [Zhang et al., 2020, Kong & Lewis, 2023, Davis et al., 2022]. A major breakthrough came with the finite-time convergence guarantees showing $\mathcal{O}(1/T)$ rates, independently developed by Yurtsever & Sra [2022] and Abbaszadehpeivasti et al. [2023]. Quite recently, Rotaru et al. [2025] provided a tighter DCA convergence analysis with applications to the convergence rates of proximal methods. We refer the interested reader to Le Thi & Pham Dinh [2018] for an in-depth discussion of the theoretical foundations, algorithmic developments, and practical applications.
Frank-Wolfe algorithms
The Frank-Wolfe algorithm [Frank & Wolfe, 1956], also known as the conditional gradient method [Levitin & Polyak, 1966], is a projection-free optimization method for constrained problems that has recently regained significant interest. This renewed attention is largely due to its suitability for problems with complicated feasible regions, where projection-based methods are often impractical. Originally designed for smooth convex optimization over polyhedral sets, the algorithm was later generalized by Levitin & Polyak [1966] to arbitrary compact convex constraint sets. The main appeal of Frank-Wolfe lies in its computational simplicity: it replaces potentially expensive or intractable projection steps with linear optimization over the constraint set. The modern revival of Frank-Wolfe methods can be traced to the influential work of Jaggi [2013], which sparked a wave of follow-up research. Since then, the Frank-Wolfe algorithm has been extended in numerous directions, including nonconvex settings [Lacoste-Julien, 2016], stochastic and block-coordinate variants [Osokin et al., 2016], and various acceleration strategies such as away steps [Guélat & Marcotte, 1986], pairwise steps [Lacoste-Julien & Jaggi, 2015], in-face directions [Freund et al., 2017], lazification [Braun et al., 2017, 2019c], blending [Braun et al., 2019b, Combettes & Pokutta, 2019, Tsuji et al., 2022], sliding techniques [Lan & Zhou, 2016, Lan et al., 2017, Qu et al., 2018], and Nesterov-style acceleration [Diakonikolas et al., 2020, Carderera et al., 2021]. For a comprehensive overview of these developments, we refer the interested reader to Braun et al. [2025].
Frank-Wolfe for DC problems
The connection between DC programming and Frank-Wolfe algorithms go back to at least Khamaru & Wainwright [2019]. Building on this, Yurtsever & Sra [2022] provided a strong relationship between DCA and Frank-Wolfe optimization via epigraphic reformulations. More recently, research has focused on developing Frank-Wolfe methods tailored to difference-of-convex optimization. For example, Millán et al. [2023] proposed a direct Frank-Wolfe approach for DC problems, showing that all accumulation points of their algorithm are stationary points of the original problem. Furthermore, under additional weak-star convexity assumptions, they established $\mathcal{O}(1/T)$ convergence rates for both duality gap measures and objective function residuals. Finally, this computational study is based on Maskan et al. [2025] as we will see further below.
Contribution
In this paper, we present a computational study that builds upon the framework of Maskan et al. [2025]. By integrating their adaptive error bound with advanced Frank-Wolfe variants, most notably the Blended Pairwise Conditional Gradients (BPCG) algorithm [Tsuji et al., 2022] with warm-starting, we obtain a highly efficient and scalable projection-free approach for difference-of-convex (DC) optimization problems, allowing for the efficient solution of large-scale instances that were previously out of reach.
We demonstrate the superiority of this approach by extensive experiments on several problem classes, such as, e.g., difference-of-convex quadratics, complicated nonconvex functions, and Quadratic Assignment Problems (QAP). In all experiments, we observe substantial gains in computational efficiency, consistently outperforming baseline methods and often achieving significant reductions in LMO calls (often by orders of magnitude) and runtime speedups exceeding factors of $100$.
2 Preliminaries and Notation
In the following we will briefly recall some basic notions and facts that we will use later in the paper and we sometimes suggestively write $x_{t}\in P$ in our formulas to hint at the later use as iterates.
2.1 The Difference-of-Convex Algorithm (DCA)
We want to solve the problem
$\min_{x\in P}\Phi(x)=\min_{x\in P}f(x)-g(x),$ (2.1)
where $P$ is a compact convex set and $\Phi(x)\doteq f(x)-g(x)$ is a difference of two convex functions with $f$ being $L$-smooth and $g$ being potentially non-smooth but admitting subgradients. For simplicity of exposition we will assume that $g$ is smooth as well and write $\nabla g$ for the gradient of $g$, however our arguments apply to the general case. In particular, the notion of gap below that uses $\nabla g$ applies also to any subgradient of $g$ in the non-smooth case.
While $\Phi(x)$ is a non-convex function, we can linearize the $g$-part of $\Phi(x)$ at $x_{t}\in P$ to obtain a convex relaxation $\hat{\Phi}_{t}(x)$ which is defined as
$\hat{\Phi}_{t}(x)\doteq f(x)-g(x_{t})-\left\langle\nabla g(x_{t}),x-x_{t}\right\rangle,$ (2.2)
and which is an $L$-smooth and convex function and by the convexity of $g$ it holds
$\Phi(x)\leq\hat{\Phi}_{t}(x).$
From this inequality we obtain the notion of the difference-of-convex gap (DC gap)
$\Delta_{x_{t}}\doteq\max_{x\in P}\Phi(x_{t})-\hat{\Phi}_{t}(x)=\max_{x\in P}f(x_{t})-f(x)-\left\langle\nabla g(x_{t}),x_{t}-x\right\rangle,$ (2.3)
and it holds that
$\Delta_{x_{t}}=\max_{x\in P}\Phi(x_{t})-\hat{\Phi}_{t}(x)\leq\max_{x\in P}\Phi(x_{t})-\Phi(x),$
i.e., the DC gap is a lower bound on the primal gap; for brevity we also write $\Delta_{t}$. Moreover, if we choose $x_{t+1}=\arg\max_{x\in P}\Phi(x_{t})-\hat{\Phi}_{t}(x)$, then it holds that
$\Delta_{t}=\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})\leq\Phi(x_{t})-\Phi(x_{t+1}),$ (2.4)
i.e., the DC gap is also a lower bound on the primal progress that can be achieved. As such, as long as the DC gap is not zero, we can make progress and together with standard arguments it follows that the DC gap is 00 at $x$ if and only if $x$ is a stationary point; see e.g., Millán et al. [2023]. Moreover, from (2.4) we immediately obtain convergence of the DC gap via standard arguments: Add up (2.4) for $t=0,\dots,T$, telescope, and then divide by $T+1$ to obtain:
$\frac{1}{T+1}\sum_{t=0}^{T}\Delta_{t}\leq\frac{1}{T+1}\sum_{t=0}^{T}\left(\Phi(x_{t})-\Phi(x_{t+1})\right)=\frac{\Phi(x_{0})-\Phi(x_{T+1})}{T+1}\leq\frac{\Phi(x_{0})-\Phi(x^{*})}{T+1},$ (2.5)
where the last inequality is by using that $x^{*}$ is a solution to (2.1), so that the average of the DC gaps (and hence its minimum!) converges to 00 at a rate of $O(1/T)$. This argument is standard and can also be made inexact with additive error $\varepsilon$; see e.g., Maskan et al. [2025], see also Section 3 for a similar reasoning. The argument also immediately gives rise to the DCA algorithm [Tao et al., 1986], which is a popular method for solving (2.1) and precisely uses the recursion as defined above, i.e.,
$x_{t+1}\leftarrow\arg\min_{x\in P}\hat{\Phi}_{t}(x).$ (DCA)
We conclude this section with collecting useful properties from the arguments above.
Fact 2.1 (Useful properties).
Let $x_{t}\in P$ with $0\leq t\leq T$ be a sequence of points (e.g., the iterates of the DCA algorithm) and $x^{*}=\arg\min_{x\in P}\Phi(x)$ be a solution to (2.1). Then, it holds that
  1. 1.
    $\Phi(x)\leq\hat{\Phi}_{t}(x)$ for all $x\in P$ and $\Phi(x_{t})=\hat{\Phi}_{t}(x_{t})$ by convexity of $g$, i.e., linearizing $g$ at $x_{t}$ provides an upper bound function for $\Phi(x)$, which is exact at $x_{t}$ itself.
  2. 2.
    $\Delta_{t}=\max_{x\in P}\Phi(x_{t})-\hat{\Phi}_{t}(x)=\max_{x\in P}f(x_{t})-f(x)-\left\langle\nabla g(x_{t}),x_{t}-x\right\rangle=\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})$, where $x_{t+1}\leftarrow\arg\min_{x\in P}\hat{\Phi}_{t}(x)$, i.e., the DCA algorithm picks $x_{t+1}$ to be the point that minimizes the linearization of $\Phi(x)$ at $x_{t}$ and hence realizes the DC gap.
  3. 3.
    $\Delta_{t}\leq\max_{x\in P}\Phi(x_{t})-\Phi(x)$, i.e., the DC gap is a lower bound for the primal gap.
2.2 The Frank-Wolfe Algorithm (FW)
The Frank-Wolfe algorithm [Frank & Wolfe, 1956], also known as the conditional gradient method [Levitin & Polyak, 1966], is a projection-free first-order method particularly well-suited for solving convex optimization problems over compact convex sets with potentially complicated constraints. In the context of DC programming, the Frank-Wolfe algorithm emerges as a natural choice for solving the convex subproblems (2.2) that arise at each iteration of the DCA algorithm; see Section 4. We will confine ourselves to the most relevant properties of the Frank-Wolfe algorithm and its vanilla variant here. However, more advanced variants exist as we will discuss later in Section 4.1.
Algorithm 1 Frank-Wolfe Algorithm [Frank & Wolfe, 1956]
1:Initial point $x_{0}\in P$, smooth convex objective function $h$
2:Iterates $x_{1},x_{2},\ldots\in P$
3:for $t=0,1,2,\ldots$ do
4:     $v_{t}\leftarrow\arg\min_{v\in P}\langle\nabla h(x_{t}),v\rangle$Linear minimization oracle}
5:     $\gamma_{t}\leftarrow\frac{2}{t+2}$Step size (open loop)}
6:     $x_{t+1}\leftarrow(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$Convex combination}
7:end for
The (vanilla) Frank-Wolfe algorithm for minimizing a smooth convex function $h(x)$ over a compact convex set $P$ is given in Algorithm 1. For a point $x\in P$, the Frank-Wolfe gap is defined as
$G(x)\doteq\max_{v\in P}\langle\nabla h(x),x-v\rangle=\langle\nabla h(x),x-v^{FW}\rangle,$ (2.6)
where $v^{FW}\in\arg\min_{v\in P}\langle\nabla h(x),v\rangle$, which is often called a Frank-Wolfe vertex and the function that provides such a solution for a given gradient is called a Linear Minimization Oracle (LMO). The Frank-Wolfe gap has the important property that $G(x)=0$ if and only if $x$ is an optimal solution to $\min_{x\in P}h(x)$. Moreover, it holds
$h(x_{t})-h(x^{*})\leq\left\langle\nabla h(x),x-x^{*}\right\rangle\leq\max_{v\in P}\left\langle\nabla h(x),x-v\right\rangle,$
where $x^{*}\in\arg\min_{x\in P}h(x)$, the first inequality is by convexity, and the second inequality is by maximality, making the Frank-Wolfe gap a natural stopping criterion and a measure of optimality. The convergence analysis of the Frank-Wolfe algorithm relies on the smoothness of the objective function; see [Jaggi, 2013, Braun et al., 2025] for details. For an $L$-smooth convex function $h$ over a compact convex set $P$ with diameter $D=\max_{x,y\in P}\|x-y\|$, the Frank-Wolfe algorithm achieves a convergence rate of
$h(x_{t})-h(x^{*})\leq\frac{2LD^{2}}{t+2}\qquad\text{ and }\qquad\min_{0\leq t\leq T}G(x_{t})\leq\frac{4LD^{2}}{T+2}.$ (2.7)
3 An adaptive error bound for DC subproblems
As shown in Maskan et al. [2025], given a target accuracy $\varepsilon>0$, we can allow for inexact solutions of the DCA subproblem, while ensuring convergence to a first-order stationary point of problem (2.1).
Theorem 3.1 (Inexact DCA with convergence guarantee; Maskan et al. [2025]).
Suppose that the sequence $x_{t}$ is generated by an inexact-DCA algorithm designed to solve the subproblems described in (2.2) approximately, in the sense that
$\hat{\Phi}_{t}(x_{t+1})-\hat{\Phi}_{t}(x)\leq\frac{\varepsilon}{2},\quad\forall x\in P,$
with $\varepsilon>0$. Then, the following bound holds:
$\min_{0\leq t\leq T}\max_{x\in P}\left\{f(x_{t})-f(x)-\langle\nabla g(x_{t}),x_{t}-x\rangle\right\}\leq\frac{\Phi(x_{0})-\Phi(x^{*})}{T+1}+\frac{\varepsilon}{2}.$ (3.1)
While Theorem 3.1 allows for practical implementations of the DCA algorithm, it still has several drawbacks. One the one hand, we need to decide on $\varepsilon>0$ ahead of time and on the other hand, we might solve the subproblems to a much higher target accuracy then actually needed for convergence, which in turn results in potentially excessive number of calls to the linear minimization oracle; our later numerical experiments confirm that this is actually happening. In a nutshell, Theorem 3.1 establishes that solving the subproblems inexactly up to additive $\varepsilon>0$, leads to a modification of the progress lower bound from (2.4) to
$\underbrace{\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})}_{\text{DC gap at $x_{t}$}}\leq\underbrace{\Phi(x_{t})-\Phi(x_{t+1})}_{\text{primal progress}}+\varepsilon,$ (3.2)
i.e., we have an additive error of $\varepsilon$ in the primal progress. In order to address these potential drawbacks Maskan et al. [2025] devised an adaptive error criterion (similar criteria have also been used before in Thuerck et al. [2023], Pokutta [2024]) that trades the additive error of $\varepsilon$ for a multiplicative error of a factor $1/2$ in the progress, via an arithmetic-geometric crossover. The resulting convergence rate is slower by a factor of $2$, however we can remove the additive error; compare (3.1) to (3.4). In particular, this allows for solving to arbitrary precision simply by letting the algorithm “run longer”, without the need for restarts or similar. We reproduce a full proof below for completeness.
Theorem 3.2 (Adaptive error bound and convergence rate for DCA; Maskan et al. [2025]).
Let $x_{t}\in P$ with $0\leq t\leq T$ be the iterates of the DCA algorithm, where the subproblems are solved inexactly satisfying the condition:
$\hat{\Phi}_{t}(x_{t+1})\leq\Phi(x_{t})-\left\langle\nabla\hat{\Phi}_{t}(x_{t+1}),x_{t+1}-v_{t+1}\right\rangle,$ (3.3)
with $v_{t+1}\in\arg\max_{v\in P}\left\langle\nabla\hat{\Phi}_{t}(x_{t+1}),x_{t+1}-v\right\rangle$, which is precisely a Frank-Wolfe vertex. Then, it holds that
$\min_{0\leq t\leq T}\Delta_{t}$ $=\min_{0\leq t\leq T}\max_{x\in P}\Phi(x_{t})-\hat{\Phi}_{t}(x)$ (3.4)
$=\min_{0\leq t\leq T}\max_{x\in P}f(x_{t})-f(x)-\left\langle\nabla g(x_{t}),x_{t}-x\right\rangle$
$\leq 2\frac{\Phi(x_{0})-\Phi(x^{*})}{T+1},$
where $x^{*}=\arg\min_{x\in P}\Phi(x)$ and $x_{t+1}$ is the inexact solution to the subproblem, i.e., the iterates $x_{t}$ converge to a stationary point of the DC problem $\min_{x\in P}\Phi(x)$ with a convergence rate of $O(1/T)$.
Proof.
Assume that (3.3) holds for all $t\in\{0,\ldots,T\}$. Then, we have that for all $x\in P$
$\hat{\Phi}_{t}(x_{t+1})$ $\leq\Phi(x_{t})-\left\langle\nabla\hat{\Phi}_{t}(x_{t+1}),x_{t+1}-v_{t+1}\right\rangle$ (3.5)
$\leq\Phi(x_{t})-(\hat{\Phi}_{t}(x_{t+1})-\hat{\Phi}_{t}(x)),$
where the second inequality is by convexity and maximality of $v_{t+1}\in P$, i.e., $\hat{\Phi}_{t}(x_{t+1})-\hat{\Phi}_{t}(x)\leq\left\langle\nabla\hat{\Phi}_{t}(x_{t+1}),x_{t+1}-x\right\rangle\leq\left\langle\nabla\hat{\Phi}_{t}(x_{t+1}),x_{t+1}-v_{t+1}\right\rangle$ for all $x\in P$. We then rearrange (3.5) to
$\hat{\Phi}_{t}(x_{t+1})\leq\frac{1}{2}\left(\Phi(x_{t})+\hat{\Phi}_{t}(x)\right),$
and subtract $\Phi(x_{t})$ from both sides and rearrange to obtain for all $x\in P$
$\frac{1}{2}\left(\Phi(x_{t})-\hat{\Phi}_{t}(x)\right)$ $\leq\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})$ (3.6)
$\leq\Phi(x_{t})-\Phi(x_{t+1}),$
where the last inequality is by convexity. In particular, for every $t$ we have
$\frac{1}{2}\Delta_{t}\doteq\frac{1}{2}\max_{x\in P}\left(\Phi(x_{t})-\hat{\Phi}_{t}(x)\right)$ $\leq\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})$ (3.7)
$\leq\Phi(x_{t})-\Phi(x_{t+1}).$
We then sum up (3.6) over all $0\leq t\leq T$, so that for all $x\in P$
$\sum_{t=0}^{T}\frac{1}{2}\Delta_{t}\leq\sum_{t=0}^{T}(\Phi(x_{t})-\Phi(x_{t+1}))=\Phi(x_{0})-\Phi(x_{T+1}),$ (3.8)
and rearranging a final time results in
$\min_{0\leq t\leq T}\Delta_{t}\leq\frac{1}{T+1}\sum_{t=0}^{T}\Delta_{t}\leq 2\frac{\Phi(x_{0})-\Phi(x_{T+1})}{T+1}\leq 2\frac{\Phi(x_{0})-\Phi(x^{*})}{T+1}.$ (3.9)
Before we continue a few remarks are in order.
Remark 3.3 (Efficiency of the subproblem error bound).
First of all, note that we can replace the factor $1/2$ and $2$ by $1-\varepsilon$ and $1/(1-\varepsilon)$, respectively via a slight modification of the argument. Furthermore, while Theorem 3.2 ensures convergence without additive errors, an obvious question is, whether we are not solving the subproblem to much higher precision compared to Theorem 3.1. This is in fact not so, to see this we can rearrange (3.3) to:
$\underbrace{\left\langle\nabla\hat{\Phi}_{t}(x_{t+1}),x_{t+1}-v_{t}\right\rangle}_{\text{Frank-Wolfe gap of $\hat{\Phi}_{t}(x)$ at $x_{t+1}$}}\leq\underbrace{\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})}_{\geq\Delta_{t}},$
i.e., we solve the subproblem $\min_{x\in P}\hat{\Phi}_{t}(x)$ until its Frank-Wolfe gap drops below $\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})$, which is at least as large as the DC gap at $x_{t}$; here $x_{t+1}$ denotes the iterate of the subsolver that it would return. In particular, for a fixed $\varepsilon>0$, as long as $\Delta_{t}\geq\varepsilon$, we do not solve the subproblems more exactly than required in Theorem 3.1. We will see later though that in fact, (3.3) is satisfied much earlier often requiring only a few iterations on the subproblem. Finally, the adaptive error bound criterion removes the need for deciding on $\varepsilon>0$ ahead of time.
4 DCA and Frank-Wolfe
In the context of our DC subproblems (2.2), applying the Frank-Wolfe algorithm is a natural choice because the subproblems $\hat{\Phi}_{t}(x)$ are $L$-smooth and convex by construction and $P$ is compact convex. Furthermore, as discussed in Remark 3.3, the Frank-Wolfe gap of the subproblem naturally appears in the adaptive error bound criterion (3.3), which allows us to easily link the required precision of the subproblem being solved with Frank-Wolfe variants to the outer DCA loop. This immediately lends itself to the algorithmic template presented in Algorithm 2, which solves the DC subproblems $\min_{x\in P}\hat{\Phi}_{t}(x)$ via some Frank-Wolfe algorithm $\mathcal{A}$ that uses as stopping criterion function a function $\tau_{t}$, i.e.,
$\langle\nabla\hat{\Phi}_{t}(y),y-v\rangle\leq\tau_{t}(y)=\Phi(x_{t})-\hat{\Phi}_{t}(y),$ (4.1)
where $v\in\arg\min_{v\in P}\langle\nabla\hat{\Phi}_{t}(y),v\rangle$ is the Frank-Wolfe vertex; we recover the variant of Theorem 3.1, simply by choosing $\tau_{t}=\varepsilon$. The algorithm can then be instantiated with different Frank-Wolfe algorithms such as vanilla Frank-Wolfe, Away-Step Frank-Wolfe, Pairwise Conditional Gradients, or Blended Pairwise Conditional Gradients for $\mathcal{A}$.
Algorithm 2 Adaptive DCA with Frank-Wolfe
1:Initial point $x_{0}\in P$, DC objective $\Phi(x)=f(x)-g(x)$, Frank-Wolfe variant $\mathcal{A}$
2:Sequence of iterates $x_{1},x_{2},\ldots\in P$
3:for $t=0,1,2,\ldots$ do
4:     Define subproblem: $\hat{\Phi}_{t}(x)=f(x)-g(x_{t})-\langle\nabla g(x_{t}),x-x_{t}\rangle$
5:     Define stopping criterion: $\tau_{t}(y)=\Phi(x_{t})-\hat{\Phi}_{t}(y)$ for $y\in P$Adaptive error bound (3.3)}
6:     $x_{t+1}\leftarrow\mathcal{A}(\hat{\Phi}_{t}(\cdot),x_{t},\tau_{t})$Call Frank-Wolfe variant}
7:end for
With $\tau_{t}=\varepsilon$ and $\mathcal{A}$ being the vanilla Frank-Wolfe algorithm (with some step-size strategy), we obtain the version in Maskan et al. [2025], who unrolled the Frank-Wolfe algorithm and baked it directly into the DCA iterations. In contrast, we kept the modular structure to vary the subsolver easily.
4.1 Leveraging advanced FW algorithms
While Maskan et al. [2025] and Millán et al. [2023] considered primarily the vanilla Frank-Wolfe algorithm [Frank & Wolfe, 1956], the former mentioning improved convergence rates in some special cases, many advanced variants of the base algorithm exist. These variants, including Away-Step FW (AFW) [Guélat & Marcotte, 1986], Fully-Corrective Frank-Wolfe (FCFW) [Holloway, 1974], Decomposition-Invariant Conditional Gradients (DICG) [Garber & Meshi, 2016, Bashiri & Zhang, 2017], Pairwise Conditional Gradients (PCG) [Lacoste-Julien & Jaggi, 2015], Blended Conditional Gradients (BCG) [Braun et al., 2019a], Boosted Frank-Wolfe (BFW) [Combettes & Pokutta, 2020], Blended Pairwise Conditional Gradients (BPCG) [Tsuji et al., 2022] often exhibit up to linear convergence rates depending on feasible region and function properties. Moreover, they provide additional benefits such as lazification [Lan et al., 2017, Braun et al., 2017, 2019c] and blending [Braun et al., 2019a, Combettes & Pokutta, 2019, Tsuji et al., 2022], where the former allows to aggressively reuse information from previous LMO calls and the latter blending in different types of descent steps for improved convergence properties. Finally, active set-based algorithms (including AFW, PCG, BCG, BFW, and BPCG), i.e., those that explicitly maintain the current iterate as convex combinations allow for warm-starting, which in turn can significantly reduce computational overhead when called repeatedly as heavily exploited in conditional gradient-based MINLP solver Boscia.jl [Hendrych et al., 2025b]. Moreover, there exist also many modern, highly-efficient line search strategies specifically for Frank-Wolfe methods [Pedregosa et al., 2020, Pokutta, 2024, Hendrych et al., 2025a] that go beyond the original open-loop step-size and the short-step strategy; we refer the reader to Braun et al. [2025] for an in-depth discussion.
Based on the benchmarks in Besançon et al. [2025], we opted for using the Blended Pairwise Conditional Gradients (BPCG) algorithm of Tsuji et al. [2022]. While it is slightly slower than the Blended Conditional Gradients algorithm of Braun et al. [2019a], it offers superior numerical stability (heavily exploited in Boscia.jl) due to blending with pairwise steps rather than projected gradient descent steps. BPCG allows for warm-starting by reusing the active set of a prior call and is adaptable to well-known function properties and properties of the feasible region. In particular, it achieves accelerated sublinear (for the sharp objectives) to linear (for strongly convex objectives) convergence rates over polytopes [Kerdreux et al., 2019, 2022] (see also Braun et al. [2025]). Similarly, BPCG achieves improved convergence rates over uniformly convex sets [Kerdreux et al., 2021] and strongly convex sets [Garber & Hazan, 2015]; also for the local case [Kerdreux et al., 2025].
We also tested other variants including Away-Step Frank-Wolfe and Pairwise Conditional Gradients, Blended Conditional Gradients, as well as the vanilla Frank-Wolfe with tracking the active set for warm-starting as these are readily available in FrankWolfe.jl. However, their performance was significantly inferior to BPCG. Moreover, we note that BPCG offers further potential improvements to reduce computational overheads, such as e.g., lazification, which potentially reduces the number of LMO calls even further, however BPCG with warm-starting and the adaptive error bound already reduces the overall number or required LMO calls so dramatically that we opted for the tighter gap control offered by the non-lazified variant. Finally, we mention that BPCG can also be combined with active set corrections from Halbey et al. [2025], basically one-shot-ing the solution of the subproblem over the current active set via an auxiliary solver. This is applicable only to the case when $\min_{x\in P}\hat{\Phi}_{t}(x)$ is a convex quadratic problem.
5 Numerical experiments
We now present our numerical experiments, demonstrating the effectiveness of the combination of the adaptive error bound criterion with warm-starting. All experiments are conducted in Julia 1.11.5 and Algorithm 2 has been implemented in the FrankWolfe.jl package [Besançon et al., 2022, 2025], so that we can directly use the latest Frank-Wolfe algorithms and line search strategies. Our DCA code will be contributed to the FrankWolfe.jl package and the code to run the computational experiments will be made separately available on GitHub. All tests were performed on a Mac Pro M1 Max. For visualization purposes, we (additively) shift the primal objective values by their minimal value to ensure non-negative values for logarithmic plotting. The early stopping via the adaptive error bound (3.3) is realized via the callback mechanism available in FrankWolfe.jl, removing the need for restarts or halving-strategies as done in Maskan et al. [2025]. All experiments are run with a single thread except for the larger Quadratic Assignment Problem (QAP) instances where FrankWolfe.jl employs the Hungarian method from the Hungarian.jl Julia package within the LMO of the Birkhoff polytope which internally is multi-threaded; this only affects absolute timing for the large QAP instances but not relative differences. Note, that still some large QAP instances require several weeks of compute time for the slower algorithms for achieving the, admittedly relatively tight, target precision.
5.1 Experimental setup
We evaluate DCA solver variants based on Algorithm 2 with different Frank-Wolfe variants and stopping criteria:
  • DCA-FW: Standard Frank-Wolfe with fixed $\varepsilon$ for subproblems as required by Theorem 3.1; basically the algorithm in Maskan et al. [2025] except for employing the much faster secant line search strategy from Hendrych et al. [2025a]. Note, that we also did preliminary tests with both the agnostic $\gamma_{t}=\frac{2}{t+2}$ step-size strategy as well as short-steps, which maximize progress over the smoothness approximation (see Dem’yanov & Rubinov [1970], Braun et al. [2025]), however both were strictly inferior to the secant line search strategy.
  • DCA-FW-ES: Standard Frank-Wolfe with adaptive stopping criterion (3.3) as required by Theorem 3.2 for early stopping.
  • DCA-BPCG: Blended Pairwise Conditional Gradients of Tsuji et al. [2022] with fixed $\varepsilon$ for subproblems as required by Theorem 3.1.
  • DCA-BPCG-ES: Blended Pairwise Conditional Gradients with adaptive stopping criterion (3.3) as required by Theorem 3.2 for early stopping.
  • DCA-BPCG-WS: Blended Pairwise Conditional Gradients with warm-starting and fixed $\varepsilon$ for subproblems as required by Theorem 3.1.
  • DCA-BPCG-WS-ES: Blended Pairwise Conditional Gradients with warm-starting and adaptive stopping criterion (3.3) as required by Theorem 3.2 for early stopping.
All algorithms use the secant line search strategy from Hendrych et al. [2025a], which is currently the fastest line search strategy for Frank-Wolfe algorithms and the default strategy in FrankWolfe.jl. In particular for quadratics this line search strategy terminates with as an optimal solution with two function evaluations. For all experiments, we impose a maximum of number of DCA iterations (between $200$ and $500$ depending on the experiment) and we stop with a DCA gap less than $10^{-6}$, which may or may not be achieved within the DCA iteration limit. We also impose an iteration limit on the inner iterations of the Frank-Wolfe variant (between $10000$ and $50000$ depending on the experiment) and we stop with a Frank-Wolfe gap less than $210^{-6}/$ 2 or when the adaptive error bound is satisfied, for those variants using it. The feasible regions vary by experiment but include the $k$-sparse polytope, probability simplex, $\ell_{1}$-ball, and the Birkhoff polytope as specified below; see Besançon et al. [2025, 2022] for detailed descriptions of these constraint sets. To assess scalability, we test various instance sizes ranging from small instances with $n=10$ to large instances with roughly $n=1000$. For each instance size, we generate multiple random instances unless stated otherwise. We also observed that the algorithms do not necessarily converge to the same primal solution, which is not unexpected as the considered problems are non-convex and the individual trajectories of the algorithms can differ vastly; see Section C for some individual runs.
We use performance profiles for various performance measures (DCA iterations, time, and LMO calls) to summarize the performance of the different algorithmic variants. The performance profiles are log-log plots and the convergence plots in Section C use a log scale on $y$-axis and also for the $x$-axis when reporting LMO calls. For the performance profiles an instance is considered solved if the DCA gap is less than $10^{-6}$ if not stated otherwise.
5.2 Summary of the experiments
Before considering the specific experiments, here we summarize our key findings. Note that these findings are very consistent across the experiments.
Adaptive error bound efficiency.
The adaptive error bound criterion proves to be highly efficient when used for early stopping (ES). Effectively it dynamically adjusts the subproblem solution precision based on the current DC gap, avoiding unnecessary computational overhead in early iterations. This can be clearly seen in the instance plots’ number of required LMO calls for the subproblems in the early iterations; see Section C.
Warm-starting benefits.
Warm-starting (WS) reuses both the last iterate $x_{t}$ and its associated active set, which contains points returned from previous linear minimization oracle (LMO) calls which then can be reused. This strategy is particularly beneficial towards the end of the optimization process when the algorithm is close to the optimal solution, so that the subproblems change only slightly and warm-starting significantly accelerates convergence in this regime.
Iteration count trade-offs.
The ES solvers sometimes exhibit slightly slower convergence in terms of DCA iterations, which aligns with the theoretical prediction in Theorem 3.2, where we trade a factor of 2 in the convergence rate for the removal of additive error terms.
LMO call efficiency.
ES solvers require significantly fewer LMO calls in the beginning phases of optimization. However, as higher accuracy is required in later iterations, the number of LMO calls naturally increases to meet the adaptive stopping criterion. Similarly, the WS solvers require significantly fewer LMO calls in the end, as warm-starting cuts out the initial phase when subproblems are very similar.
Combination of adaptive error bound together with warm-starting.
While each of the two strategies ES or WS are very beneficial in terms of reducing computational cost, the combination of the two strategies is the most efficient by a huge margin, enabling multiple orders of magnitude fewer LMO calls and reducing running times by a similar factor as shown in Figure 1. In particular, the combination of the two strategies enables us to tackle instance sizes that would otherwise be out of reach. In fact, in many cases, the improvement is so substantial that we cannot precisely quantify it because the baseline solvers fail to complete within the time limit.
Results across all considered instances (except for QAP) are summarized in Table 1 and Figure 1. The table uses shifted geometric means as customary; more results and examples of representative individual runs are to be found in Section C. We excluded QAP from this overview as the instances typically do not solve to optimality and we thus we report these separately in detail in Section 5.5. As can be seen, the DCA-BPCG-WS-ES variant (combining warm-starting and early stopping) is the most efficient both in time and LMO calls being often by multiple orders of magnitude faster (geometric means) than the other variants. For a very small number of experiments ($\approx 1\%$) it seems that DCA-BPCG-WS outperform DCA-BPCG-WS-ES as can be seen that the very top of the performance profile chart. However this is mostly an artifact of our experimental design and the hard cap on DCA iterations. By Theorem 3.2 we know that the primal progress may be slower in DCA iterations when using the adaptive error bound criterion while each of the DCA iterations is much faster. That is precisely what we see here: these select few experiments finish via DCA-BPCG-WS very close to the DCA iteration limit and the DCA-BPCG-WS-ES variant would require a few extra iterations to close the gap. This can be also seen when compared to the required compute time: DCA-BPCG-WS-ES requires several orders of magnitude less time (geometric mean) than DCA-BPCG-WS so that a few extra iterations are not a problem.
Figure 1: Performance profile over all tested instances (excluding QAP) as described in Section 5.1. The plot demonstrates the superior efficiency of the DCA-BPCG-WS-ES variant (combining warm-starting and early stopping) across the entire range of problem instances.
Table 1: Performance Statistics over all instances (excluding QAP; see Section 5.5) [Shifted Geometric Means, shift: $1.0$].
Size DCA-FW DCA-BPCG DCA-BPCG-WS DCA-FW-ES DCA-BPCG-ES DCA-BPCG-WS-ES
Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO
$n=10$ (10) 24.65 1.57 6152.00 8.90 0.06 106.19 8.90 0.04 114.97 20.05 0.76 434.32 16.44 0.85 483.16 9.65 3.04e-03 42.05
$n=20$ (10) 63.91 5.20 4.98e+05 16.06 0.16 3361.08 16.24 0.16 1544.55 71.32 4.87 7.82e+04 71.15 5.20 6.41e+04 16.50 2.83e-03 90.21
$n=30$ (10) 61.92 4.32 1.06e+05 21.09 0.28 2861.19 21.26 0.37 2673.25 29.04 0.81 1056.74 28.95 1.14 705.45 19.64 6.38e-03 154.21
$n=50$ (10) 67.22 8.51 8.19e+04 34.84 2.15 1.36e+04 22.69 0.75 3501.54 50.46 5.15 7140.91 49.67 4.52 5760.02 30.31 0.07 330.21
$n=100$ (20) 144.57 29.09 8.85e+05 44.70 1.89 3.83e+04 41.49 1.15 6279.90 118.36 15.92 1.22e+05 125.41 17.24 1.24e+05 36.18 0.09 284.85
$n=150$ (10) 60.89 28.57 1.13e+05 39.72 7.74 3.90e+04 26.36 2.94 6116.85 54.70 13.09 1.05e+04 50.80 11.65 9254.55 27.58 0.02 135.35
$n=200$ (20) 155.52 78.73 1.39e+06 81.08 12.64 3.01e+05 58.71 3.38 5.00e+04 175.92 42.56 3.38e+05 165.38 34.04 2.33e+05 64.16 0.13 711.14
$n=300$ (20) 221.76 212.14 2.16e+06 123.29 37.87 6.14e+05 81.22 12.88 2.24e+05 246.98 123.15 9.43e+05 261.51 122.23 9.53e+05 90.48 0.57 2106.04
$n=400$ (20) 259.76 292.36 1.74e+06 218.06 93.76 1.05e+06 91.87 15.53 1.36e+05 279.67 228.21 1.54e+06 260.28 146.31 9.45e+05 80.69 0.63 1876.60
$n=500$ (20) 180.95 291.42 1.14e+06 131.41 79.50 5.66e+05 70.59 21.53 1.68e+05 223.16 180.22 7.62e+05 206.73 132.95 5.74e+05 87.64 2.34 8266.83
$n=600$ (10) 400.69 408.89 2.33e+06 217.09 135.01 1.70e+06 121.52 30.45 3.78e+05 399.20 306.11 1.86e+06 402.62 222.93 1.39e+06 135.39 1.35 7345.49
$n=700$ (10) 421.90 298.50 1.92e+06 298.63 109.60 9.26e+05 92.27 12.81 9.15e+04 399.95 260.05 1.74e+06 306.37 127.02 9.67e+05 80.22 1.14 3409.07
$n=800$ (10) 365.52 579.95 3.42e+06 356.54 346.69 3.09e+06 128.18 40.42 3.97e+05 368.72 251.60 1.38e+06 360.66 209.15 1.35e+06 125.44 1.66 5473.48
$n=900$ (10) 221.82 331.17 8.10e+05 159.99 191.99 6.21e+05 88.35 46.95 1.90e+05 200.90 121.79 2.68e+05 170.29 87.36 2.10e+05 86.70 2.21 5270.98
$n=1000$ (10) 236.52 239.21 4.59e+05 151.12 90.62 2.32e+05 88.63 27.89 7.70e+04 228.92 204.13 3.94e+05 197.63 129.56 2.79e+05 88.62 1.92 3084.11
5.3 Difference of Convex Quadratic Functions
For this straightforward experiment, we generate instances where both $f$ and $g$ are convex quadratic functions of the form
$f(x)=\frac{1}{2}x^{T}Ax+a^{T}x+c\quad\text{and}\quad g(x)=\frac{1}{2}x^{T}Bx+b^{T}x+d,$ (5.1)
where $A,B\in\mathbb{R}^{n\times n}$ are positive definite matrices (constructed as $M^{T}M+0.1I$ for random matrices $M$), $a,b\in\mathbb{R}^{n}$ are random vectors, and $c,d\in\mathbb{R}$ are random scalars. The feasible region is the probability simplex $P=\{x\in\mathbb{R}^{n}:x\geq 0,\sum_{i=1}^{n}x_{i}=1\}$. We ran two separate tests, once with smaller to moderate instance sizes (see Table 2 and Figure 2), where we imposed an inner iteration limit of $10000$ iterations and a limit of $200$ DCA iterations, and once with larger instance sizes (see Table 3 and Figure 3 and Figure 3), where we imposed an inner iteration limit of $50000$ iterations and a limit of $500$ DCA iterations. In both cases, we set the DCA gap tolerance to $10^{-6}$ and the Frank-Wolfe gap tolerance to $210^{-6}/$ 2.
As can be seen from the results, in both tests BPCG-WS-ES, combining warm-starting and early stopping is the most efficient outperforming any other variant by a large margin.
Figure 2: Performance profile over small to medium-sized instances of differences of convex quadratics from Section 5.3.
Table 2: Performance Statistics over small to medium-sized instances of differences of convex quadratics from Section 5.3 [Shifted Geometric Means, shift: $1.0$].
Size DCA-FW DCA-BPCG DCA-BPCG-WS DCA-FW-ES DCA-BPCG-ES DCA-BPCG-WS-ES
Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO
$n=10$ (5) 6.36 0.20 92.98 4.70 0.05 19.56 4.70 2.18e-03 18.88 4.70 0.02 16.41 4.70 0.04 16.41 4.91 2.48e-03 17.17
$n=20$ (5) 75.56 3.84 6.39e+05 15.25 8.36e-03 886.08 15.25 2.02e-03 203.12 95.02 4.12 1.83e+05 95.02 4.47 1.80e+05 15.78 1.53e-03 79.17
$n=30$ (5) 57.96 1.72 3.26e+04 21.80 5.06e-03 463.72 22.00 2.18e-03 197.33 35.67 0.58 1611.83 22.70 2.25e-03 150.16 20.12 1.78e-03 116.94
$n=50$ (5) 63.54 1.83 1.02e+04 20.82 0.03 716.17 20.59 2.75e-03 186.30 23.52 5.20e-03 246.55 22.78 5.74e-03 175.84 21.08 2.32e-03 121.46
$n=100$ (5) 64.10 9.01 5.49e+05 24.93 0.45 1.56e+04 25.03 0.09 1173.44 62.98 5.70 2.82e+04 62.98 6.11 2.77e+04 22.97 5.29e-03 116.52
$n=150$ (5) 26.03 2.67 9271.30 22.06 1.42 3921.14 14.07 0.09 221.64 22.06 1.07 395.61 21.74 1.05 384.23 13.74 3.01e-03 58.29
$n=200$ (5) 131.48 44.42 1.26e+06 112.80 20.27 7.01e+05 58.81 1.35 1.39e+04 132.52 22.72 2.76e+05 132.52 23.19 2.61e+05 54.49 0.02 355.98
$n=300$ (5) 137.30 97.19 1.32e+06 96.82 19.64 3.70e+05 82.72 7.89 1.92e+05 139.04 31.12 2.62e+05 140.70 31.71 2.67e+05 82.72 0.05 750.10
$n=400$ (5) 200.00 146.02 1.98e+06 182.94 50.85 1.07e+06 72.22 6.31 1.16e+05 200.00 96.07 1.56e+06 200.00 91.76 1.54e+06 72.24 0.28 944.28
$n=500$ (5) 51.26 37.41 9.04e+04 38.88 17.77 5.40e+04 23.73 4.71 1.62e+04 69.20 20.56 5.52e+04 78.94 37.88 1.06e+05 37.41 0.40 865.13
Figure 3: Performance profile over medium to large-sized instances of differences of convex quadratics from Section 5.3.
Table 3: Performance Statistics over medium to large-sized instances of differences of convex quadratics from Section 5.3 [Shifted Geometric Means, shift: $1.0$].
Size DCA-FW DCA-BPCG DCA-BPCG-WS DCA-FW-ES DCA-BPCG-ES DCA-BPCG-WS-ES
Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO
$n=100$ (10) 231.25 23.32 9.58e+05 47.31 0.47 2.03e+04 47.17 0.11 2362.48 124.64 7.56 7.04e+04 169.42 12.94 1.74e+05 37.85 6.84e-03 230.11
$n=200$ (10) 295.57 85.48 2.54e+06 108.16 8.94 2.81e+05 76.29 2.16 5.71e+04 294.68 61.78 1.06e+06 297.27 68.48 1.13e+06 84.33 0.08 981.80
$n=300$ (10) 360.18 188.71 3.51e+06 144.27 27.36 6.37e+05 73.15 5.60 1.42e+05 365.52 138.48 1.84e+06 500.00 240.35 4.61e+06 81.69 0.25 1985.46
$n=400$ (10) 405.53 218.76 1.85e+06 337.29 116.75 1.38e+06 95.30 9.11 7.60e+04 390.92 175.82 1.65e+06 390.92 176.12 1.67e+06 94.94 0.71 3384.47
$n=500$ (10) 398.44 402.08 3.92e+06 272.77 142.85 2.23e+06 101.70 20.20 3.62e+05 500.00 403.29 4.47e+06 406.82 224.01 2.54e+06 101.19 1.06 1.17e+04
$n=600$ (10) 400.69 408.89 2.33e+06 217.09 135.01 1.70e+06 121.52 30.45 3.78e+05 399.20 306.11 1.86e+06 402.62 222.93 1.39e+06 135.39 1.35 7345.49
$n=700$ (10) 421.90 298.50 1.92e+06 298.63 109.60 9.26e+05 92.27 12.81 9.15e+04 399.95 260.05 1.74e+06 306.37 127.02 9.67e+05 80.22 1.14 3409.07
$n=800$ (10) 365.52 579.95 3.42e+06 356.54 346.69 3.09e+06 128.18 40.42 3.97e+05 368.72 251.60 1.38e+06 360.66 209.15 1.35e+06 125.44 1.66 5473.48
$n=900$ (10) 221.82 331.17 8.10e+05 159.99 191.99 6.21e+05 88.35 46.95 1.90e+05 200.90 121.79 2.68e+05 170.29 87.36 2.10e+05 86.70 2.21 5270.98
$n=1000$ (10) 236.52 239.21 4.59e+05 151.12 90.62 2.32e+05 88.63 27.89 7.70e+04 228.92 204.13 3.94e+05 197.63 129.56 2.79e+05 88.62 1.92 3084.11
5.4 Hard DC function example
As a more challenging example we consider a DC optimization problem that combines quadratic, exponential, and logistic components. Specifically, we minimize $\Phi(x)=f(x)-g(x)$ where
$f(x)=\frac{1}{2}x^{T}Ax+a^{T}x+\frac{1}{n}\exp\left(\frac{1}{n}\sum_{i=1}^{n}c_{i}x_{i}\right),$
and
$g(x)=\frac{1}{2}x^{T}Bx+b^{T}x+0.1\sum_{i=1}^{n}\log(1+\exp(d_{i}x_{i})),$
over the $k$-sparse polytope $C_{\tau,k}=B_{1}(k\tau)\cap B_{\infty}(\tau)=\{x\in\mathbb{R}^{n}:\|x\|_{1}\leq k\tau,\|x\|_{\infty}\leq\tau\}$ with $\tau=10$ and $k=10$, which is the intersection of the $\ell_{1}$-ball of radius $k\tau$ and the $\ell_{\infty}$-ball of radius $\tau$. This example is particularly challenging because the exponential term in $f(x)$ can lead to rapid growth, while the logistic terms in $g(x)$ introduce additional nonlinearity. The $k$-sparse polytope further complicates the optimization landscape. The matrices $A$, $B$ and the vectors $a,b,c,d\in\mathbb{R}^{n}$ are constructed as in Section 5.3.
Results are reported in Table 4 and Figure 4 and match the observations from Section 5.3.
Figure 4: Performance profile over hard DC instances from Section 5.4.
Table 4: Performance Statistics for hard DC instances from Section 5.4 [Shifted Geometric Means, shift: $1.0$].
Size DCA-FW DCA-BPCG DCA-BPCG-WS DCA-FW-ES DCA-BPCG-ES DCA-BPCG-WS-ES
Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO
$n=10$ (5) 88.42 4.53 4.03e+05 16.21 0.06 557.81 16.21 0.07 675.40 76.80 2.03 1.09e+04 52.38 2.28 1.35e+04 18.21 3.60e-03 101.01
$n=20$ (5) 54.03 6.94 3.88e+05 16.91 0.34 1.27e+04 17.30 0.35 1.17e+04 53.47 5.72 3.33e+04 53.21 6.03 2.29e+04 17.26 4.14e-03 102.78
$n=30$ (5) 66.14 9.44 3.41e+05 20.41 0.63 1.76e+04 20.54 0.89 3.61e+04 23.61 1.07 692.70 36.84 3.58 3300.69 19.17 0.01 203.28
$n=50$ (5) 71.10 30.98 6.57e+05 57.88 8.69 2.57e+05 25.00 2.05 6.55e+04 107.00 36.66 2.06e+05 107.00 29.27 1.88e+05 43.39 0.15 894.76
$n=100$ (5) 126.88 137.60 1.22e+06 71.12 21.09 3.36e+05 53.00 14.78 2.37e+05 200.00 166.19 1.60e+06 136.40 79.22 2.86e+05 51.84 0.40 1062.69
$n=150$ (5) 140.70 237.07 1.39e+06 70.91 30.58 3.87e+05 48.67 13.28 1.68e+05 133.53 95.10 2.80e+05 116.97 77.18 2.22e+05 54.40 0.04 312.59
$n=200$ (5) 50.51 117.98 4.58e+05 32.48 15.49 1.49e+05 34.58 14.64 1.38e+05 82.93 37.51 4.22e+04 63.52 11.91 8853.26 43.61 0.36 744.93
$n=300$ (5) 135.47 583.05 1.33e+06 114.63 136.54 9.50e+05 98.29 94.91 6.53e+05 200.00 379.09 8.90e+05 132.52 120.06 1.45e+05 121.35 2.73 6649.18
$n=400$ (5) 138.18 1042.07 1.36e+06 108.40 111.16 5.99e+05 108.57 98.88 5.18e+05 200.00 908.32 1.30e+06 149.94 160.82 1.85e+05 65.05 0.91 1146.13
$n=500$ (5) 130.43 1170.53 1.20e+06 101.83 107.13 3.83e+05 99.72 99.51 3.77e+05 142.28 305.00 3.06e+05 139.04 162.55 1.59e+05 152.95 20.07 3.96e+04
5.5 Quadratic Assignment Problem
In this test we solve a DC relaxation of the Quadratic Assignment Problem (QAP). The QAP aims to find an optimal alignment between two matrices $A,B\in\mathbb{R}^{n\times n}$, and it is well known that this is equivalent to minimizing a quadratic objective function over the set of permutation matrices
$\underset{X\in\mathbb{R}^{n\times n}}{\min}\left\langle A^{\top}X,XB\right\rangle\quad\mathrm{s.t.}\quad X\in\{0,1\}^{n\times n},~{}X1_{n}=X^{\top}1_{n}=1_{n},$ (5.2)
where $1_{n}$ denotes the $n$-dimensional vector of ones. As this combinatorial optimization problem is NP-hard, we want to solve it over the linear relaxation of the constraints, where we drop the integrality required:
$\underset{X\in\mathbb{R}^{n\times n}}{\min}~{}~{}\left\langle A^{\top}X,XB\right\rangle\quad\mathrm{s.t.}\quad X\in[0,1]^{n\times n},~{}X1_{n}=X^{\top}1_{n}=1_{n},$ (5.3)
i.e., we want to minimize a nonconvex quadratic over the Birkhoff polytope $\mathcal{B}_{n}=\{X\in\mathbb{R}^{n\times n}:X1_{n}=1_{n},X^{\top}1_{n}=1_{n},X\geq 0\}$. This solution can then often be used for rounding to obtain an actual feasible assignment of the original problem; see, e.g., Vogelstein et al. [2015], see also Maskan et al. [2025] for an in-depth discussion.
In order to apply DCA, we need to reformulate the QAP and in fact decompose the objective $\left\langle A^{\top}X,XB\right\rangle$ into a difference of convex functions. Several decompositions are possible and we opted for the approach in Maskan et al. [2025] based on their tests and for comparability:
$f(X)=\frac{1}{4}\|A^{\top}X+XB\|_{F}^{2},$
and
$g(X)=\frac{1}{4}\|A^{\top}X-XB\|_{F}^{2},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm. This gives us $\left\langle A^{\top}X,XB\right\rangle=f(X)-g(X)$ since
$f(X)-g(X)=\frac{1}{4}\left(\|A^{\top}X+XB\|_{F}^{2}-\|A^{\top}X-XB\|_{F}^{2}\right)=\left\langle A^{\top}X,XB\right\rangle.$
Note that both $f$ and $g$ are convex functions as compositions of convex quadratic forms with linear transformations.
While projection onto the Birkhoff polytope is computationally expensive, the LMO can be implemented efficiently using the Hungarian algorithm, which is readily available in the FrankWolfe.jl package. We evaluate the algorithms on instances from QAPLIB [Burkard et al., 1997], a well-established benchmark library for quadratic assignment problems111available at https://coral.ise.lehigh.edu/data-sets/qaplib/qaplib-problem-instances-and-solutions/. QAPLIB contains a wide range of QAP instances that have been used extensively in the literature to benchmark QAP algorithms. From the 136 instances in QAPLIB, we successfully processed 97 instances (71.3%) for our experiments. The remaining 39 instances (28.7%) could not be processed due to various data format inconsistencies or parsing issues; for details see Section A.
For our experiments, we used a DCA iteration limit of $500$ and an inner iteration limit for the Frank-Wolfe iterations of $50000$. The instance characteristics can be found in Table 5. Note that the actual problem dimension is quadratic in $n$. Due to their considerable size and the substantially higher LMO complexity (as we have to now run the Hungarian algorithm compared to e.g., sorting before), these instances are much harder than the previous examples, also often coming with objective values that are orders of magnitude larger, so that the additive tolerance is relatively tighter. Thus, the target DCA tolerance of $10^{-6}$ could not always be achieved, even for the fastest algorithm configuration.
Given these computational challenges, we present two complementary performance profiles for these experiments: the classical performance profile in Figure 5, and a modified performance profile in Figure 6 where we consider all instances to be “solved” at their final iteration. This second profile allows us to compare the computational cost—both in terms of runtime and LMO calls—required to reach the DCA iteration limit across different algorithm variants.
Also note that in Table 5, the results seem to be a little bit more noisy, which in fact is due to often having only one QAP instance in a given size bracket. Due to nonconvexity, the optimization paths can differ signficantly, so that sometimes the slower algorithms can be faster as they converged to a different stationary point.
Figure 5: Performance profile over Quadratic Assignment Problem instances from Section 5.5.
Figure 6: Modified performance profile over Quadratic Assignment Problem instances from Section 5.5, where we consider all instances to be solved in their last iteration.
Table 5: Performance Statistics Quadratic Assignment Problem instances from Section 5.5 [Shifted Geometric Means, shift: $1.0$].
Size DCA-FW DCA-BPCG DCA-BPCG-WS DCA-FW-ES DCA-BPCG-ES DCA-BPCG-WS-ES
Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO Iter Time LMO
$n=10$ (4) 500.00 120.53 5.00e+06 307.47 70.15 2.96e+06 177.98 31.30 1.43e+06 500.00 88.75 3.52e+06 500.00 105.81 3.76e+06 168.84 1.85 7.42e+04
$n=12$ (18) 500.00 193.50 5.00e+06 500.00 182.21 4.96e+06 164.44 36.37 9.97e+05 500.00 149.91 3.71e+06 500.00 159.10 3.71e+06 161.41 3.83 8.45e+04
$n=14$ (4) 500.00 286.42 5.00e+06 500.00 258.41 4.95e+06 96.73 12.54 2.48e+05 500.00 270.04 4.57e+06 500.00 287.84 4.63e+06 101.40 2.79 4.82e+04
$n=15$ (16) 500.00 315.63 5.00e+06 474.88 280.30 4.71e+06 201.66 79.17 1.37e+06 500.00 246.40 3.75e+06 500.00 230.92 3.30e+06 241.32 12.58 1.51e+05
$n=16$ (13) 336.96 243.49 1.77e+06 280.60 195.65 1.44e+06 58.17 12.45 9.34e+04 326.58 223.25 1.53e+06 326.58 242.75 1.55e+06 52.80 1.25 8659.07
$n=17$ (2) 500.00 402.47 5.00e+06 500.00 334.74 4.89e+06 80.39 30.59 4.50e+05 500.00 376.93 4.64e+06 500.00 394.03 4.56e+06 88.36 4.96 6.14e+04
$n=18$ (4) 500.00 457.20 5.00e+06 334.71 267.48 3.22e+06 141.19 41.64 5.01e+05 500.00 444.26 4.61e+06 500.00 460.44 4.56e+06 216.88 10.29 1.16e+05
$n=20$ (8) 500.00 629.25 5.00e+06 500.00 574.35 4.97e+06 228.20 148.40 1.35e+06 500.00 471.76 3.66e+06 500.00 543.62 3.96e+06 219.57 11.92 9.83e+04
$n=21$ (1) 500.00 665.66 5.00e+06 500.00 599.23 4.92e+06 300.00 109.76 9.54e+05 500.00 539.76 3.77e+06 500.00 545.62 3.61e+06 455.00 49.25 4.09e+05
$n=22$ (3) 500.00 695.69 5.00e+06 500.00 620.95 4.95e+06 298.58 179.39 1.53e+06 500.00 570.54 4.02e+06 500.00 612.17 4.19e+06 411.68 51.99 4.19e+05
$n=24$ (1) 500.00 865.14 5.00e+06 500.00 758.07 4.93e+06 432.00 180.65 1.24e+06 500.00 704.49 3.95e+06 500.00 721.46 3.86e+06 101.00 7.98 4.99e+04
$n=25$ (4) 500.00 1096.12 5.00e+06 500.00 957.49 4.93e+06 350.22 458.78 2.55e+06 500.00 832.00 3.76e+06 500.00 939.65 4.05e+06 370.15 76.56 4.22e+05
$n=26$ (6) 500.00 1472.84 5.00e+06 500.00 1522.93 5.00e+06 500.00 1262.66 4.93e+06 500.00 1097.81 3.55e+06 500.00 1193.52 3.63e+06 500.00 163.96 6.14e+05
$n=27$ (1) 500.00 1109.44 5.00e+06 500.00 983.31 4.96e+06 202.00 205.24 1.06e+06 500.00 927.43 3.87e+06 500.00 1106.67 4.64e+06 196.00 39.40 1.88e+05
$n=28$ (1) 500.00 1267.68 5.00e+06 500.00 1076.75 4.95e+06 236.00 137.59 6.31e+05 500.00 853.08 3.33e+06 500.00 749.71 2.75e+06 346.00 20.35 9.48e+04
$n=30$ (3) 500.00 1931.15 5.00e+06 500.00 1788.34 4.96e+06 463.39 1144.63 3.56e+06 500.00 1485.99 3.91e+06 500.00 1518.37 3.86e+06 377.48 107.08 3.33e+05
$n=32$ (8) 500.00 1753.54 5.00e+06 500.00 1683.56 5.00e+06 149.63 459.75 1.44e+06 500.00 1606.51 4.43e+06 500.00 1649.78 4.32e+06 129.19 91.05 2.63e+05
$n=35$ (2) 500.00 2507.13 5.00e+06 500.00 2208.98 4.96e+06 324.90 1310.80 2.98e+06 500.00 1680.52 3.47e+06 500.00 2370.69 4.43e+06 500.00 81.36 1.73e+05
$n=40$ (3) 500.00 3294.19 5.00e+06 500.00 2908.98 4.97e+06 365.42 1622.86 2.89e+06 500.00 2707.80 4.01e+06 500.00 2625.10 3.96e+06 341.45 206.14 3.72e+05
$n=50$ (3) 500.00 5879.99 5.00e+06 500.00 5006.42 4.96e+06 500.00 3038.16 3.15e+06 500.00 4199.93 3.73e+06 500.00 3423.85 2.82e+06 500.00 341.62 3.50e+05
$n=60$ (2) 500.00 8838.29 5.00e+06 500.00 7606.62 4.96e+06 500.00 6400.82 4.31e+06 500.00 6513.19 4.01e+06 500.00 6672.89 3.61e+06 417.15 1106.33 7.34e+05
$n=64$ (2) 500.00 1.48e+04 5.00e+06 500.00 1.47e+04 5.00e+06 334.75 8705.74 3.35e+06 500.00 1.79e+04 4.64e+06 500.00 506.80 1.08e+05 488.36 1.93e+04 4.36e+06
$n=80$ (1) 500.00 1.05e+04 5.00e+06 500.00 8168.45 4.93e+06 345.00 4330.87 2.56e+06 500.00 1611.22 7.84e+05 500.00 3092.43 1.42e+06 500.00 1161.06 7.54e+05
$n=100$ (2) 500.00 9267.98 1.58e+06 500.00 8585.76 1.58e+06 500.00 7975.16 1.48e+06 500.00 4438.56 6.72e+05 500.00 3856.15 6.33e+05 500.00 542.98 1.07e+05
$n=128$ (1) 500.00 7165.42 5.02e+05 500.00 8143.55 5.02e+05 400.00 5057.93 4.01e+05 500.00 8224.50 4.88e+05 500.00 9925.79 4.85e+05 500.00 7541.98 4.81e+05
$n=150$ (1) 500.00 1.49e+04 5.02e+05 500.00 1.72e+04 5.02e+05 500.00 1.69e+04 5.02e+05 500.00 1.47e+04 4.80e+05 500.00 1.52e+04 4.45e+05 500.00 1.57e+04 4.53e+05
Acknowledgements
Research reported in this paper was partially supported through the Research Campus Modal funded by the German Federal Ministry of Education and Research (fund numbers 05M14ZAM,05M20ZBM) and the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+.
B Boosted DCA
In this section we present the preliminary results of the boosted DCA variant. We implemented a simple two level line search via refined grids to compute the step size; relatively exact but not the most efficient. The results are presented in Figure 7. We can see that the boosted DCA variant is basically identical to the unboosted DCA-BPCG-WS-ES variant. Neither is there a significant improvement in DCA iterations nor is it much slower in time due to the additional line search. Note however, that these are preliminary results and an in-depth analysis is left for future work.
Figure 7: Performance profile over difference of convex quadratic functions as initial test of the effect of boosted DCA. The boosted DCA variant is denoted by DCA-BPCG-WS-ES-BT and we can see that its performance is basically identical to the unboosted DCA-BPCG-WS-ES variant.
C Additional Numerical Experiments
In the following we present some selected runs to complement the summary statistics from before with some more details.
C.1 Difference of Convex Quadratic Functions
In this section we present runs of some selected Difference of Convex Quadratic Functions examples from Section 5.3 in Figures 8, 9, 10 and 11; we took the largest sizes for presentation here as the results are more interesting.
C.2 Hard DC function example
In this section we present runs of the hard DC function example from Section 5.4 in Figures 12, 13, 14 and 15.
C.3 Quadratic Assignment Problem
In this section we present runs of some selected QAP instances from Section A in Figures 16, 17, 18 and 19.
Figure 8: Run of the different variants on a Difference of Convex Quadratics example with seed $a4fd7a7efdf1f7d1$ with size $n=1000$ from Section 5.3.
Figure 9: Run of the different variants on a Difference of Convex Quadratics example with seed $e337c88623524354$ with size $n=1000$ from Section 5.3.
Figure 10: Run of the different variants on a Difference of Convex Quadratics example with seed $f308059bfbdf1ef$ with size $n=1000$ from Section 5.3.
Figure 11: Run of the different variants on a Difference of Convex Quadratics example with seed $5931de02be2505f$ 593 with size $n=500$ from Section 5.3.
Figure 12: Run of the different variants on the hard DC function example with size $n=400$ from Section 5.4.
Figure 13: Run of the different variants on the hard DC function example with size $n=500$ from Section 5.4.
Figure 14: Run of the different variants on the hard DC function example with size $n=300$ from Section 5.4.
Figure 15: Run of the different variants on the hard DC function example with size $n=150$ from Section 5.4.
Figure 16: Run of the different variants on the small-sizedesc32b QAP instance from Section A.
Figure 17: Run of the different variants on the medium-sized tai60b QAP instance from Section A.
Figure 18: Run of the different variants on the medium-sized tai64c QAP instance from Section A. For this instance the DCA gaps are oscillating wildly.
Figure 19: Run of the different variants on the large-sized tai100a QAP instance from Section A.